A condensed balance sheet for Bradford Corporation prepared at the end of the year appears as follows.

Assets Liabilities & Stockholders' Equity
Cash $ 107,000 Notes payable (due in 6 months) $ 44,000
Accounts receivable 120,000 Accounts payable 118,000
Inventory 270,000 Long-term liabilities 328,000
Prepaid expenses 60,000 Capital stock, $5 par 300,000
Plant & equipment (net) 570,000 Retained earnings 427,000
Other assets 90,000
Total $ 1,217,000 Total $ 1,217,000

During the year, the company earned a gross profit of $1,116,000 on sales of $2,950,000. Accounts receivable, inventory, and plant assets remained almost constant in amount throughout the year, so year-end figures may be used rather than averages.

a. Compute the current ratio. (Round your answer to 2 decimal place.)

b. Compute the quick ratio. (Round your answer to 2 decimal place.)

c. Compute the working capital.

d. Compute the debt ratio. (Round your percentage answers to nearest whole percent. i.e. 0.1234 as 12%.)

e. Compute the accounts receivable turnover (all sales were on credit). (Round your answer to 2 decimal places.)

f. Compute the inventory turnover. (Round your answer to 2 decimal places.)

g. Compute the book value per share of capital stock. (Round your answer to 2 decimal places.)
Answer #1
assets liabilities
cash 107000 note payable 44000
account receivable 120000 account payable 118000
inventory 270000 current liability 162000
prepaid expense 60000 long term liability 328000
Total liability 490000 162000+328000
total current asset 557000 capital stock 300000
plant and equipment 570000 retained earnings 427000
other asset 90000
Total asset 1217000 total liability 1217000

a. current ratio = current asset/ current liability

=557000/162000

=3.44

b.quick ratio

cash + account receivable/ current liability

(107000+120000)/162000

=1.40

c.working capital

=current asset-current liability

=557000-162000

=395000$

d.debt ratio

Total liability/ total asset

=490000/1217000

=40.26%

e.account receivable turnover

credit sales/average account receivable

2950000/120000

=24.58

f.inventory turnover

cost of goods sold/inventory

=(sales-gross profit)/inventory

(2950000-1116000)/270000

=1834000/270000

=6.79

g.book value

stockholder's equity / number of common shares

=[300000+427000]/[300000/$ 5 per share]

=727000/60000

=12.12